# Gianluca Monaci          Rafael S. Rezende          Romain Deffayet          Gabriela Csurka Guillaume Bono          Hervé Déjean          Stéphane Clinchant          Christian Wolf

### v1

@@ A.3 DINOv2-Graph
Selecting a path on this semantic graph (DINOv2-Graph) is not necessarily equivalent to a “physical path” (see examples in Fig. 6). In order to enforce selection of edges between images that have high overlap probability (i.e.represent the same part of the scene) we made two further modifications to the above graph. First, as we observed that a DINOv2 similarity score above $th=0.75$ between L2 normalized features reflects, in most cases, a good overlap between the two images, we only kept edges for which the similarity was above this threshold, removing all the others. Second, in order to enforce Dijkstra’s algorithm to select paths with high similarity (low cost), we used the weights $w_{ij}=\sqrt{(1-s_{ij})}$, where $s_{ij}$ is the DINOv2 similarity between the L2 normalized features.
During both training and inference, at each step, the agent finds in $\mathcal{D}$ the most similar images $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ to the observation $\mathbf{x}_{t}$ and goal $\mathbf{g}$, respectively, and computes the shortest path between $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ on the DINOv2-Graph. Let $\mathcal{P}=\{r_{1}^{\mathbf{x}_{t}},r_{1},r_{2},...r_{n},r_{1}^{\mathbf{g}}\}$ be the set of nodes on the shortest path and $C$ the context size. If $n+2>C$, we sample images randomly from $\mathcal{P}$ and update all images in the context. If $n+2\leq C$, we use all images in $\mathcal{P}$ to update $n+2$ images in the context. If no path was found, $n=2$ hence we only update two elements of the context with $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$.
@@
Table 10: [Uncaptioned image]  Impact of the graph structure on RANa with DINOv2-Graph - SR (SPL) for ImageNav with retrieval using different graph structures at test time. All RANa models were trained with Sparse Weighted Graph (SWG). In bolded best results the PG variant is not considered as it uses the camera poses of images in $\mathcal{D}$, we only add it for comparison.
@@
Analyzing the choice of the graph structure – We trained our models on the Gibson dataset where we had around $10$k images per scene in the retrieval dataset $\mathcal{D}$. We selected the checkpoint obtained at $130$M steps for all models. During testing we used a set of $|\mathcal{D}|=1$k images per scene. In this section we study how the graph structure affects performance. The default model, used for the results reported in Tab. 2 of the main paper, builds a weighted graph by removing edges with similarity below $th=0.75$ and uses $w_{ij}=\sqrt{(1-s_{ij})}$ as weights. We call this Sparse Weighted Graph (SWG). We evaluate three additional strategies to construct the DINOv2-Graph. (i) We consider the same graph, but instead of using $w_{ij}=\sqrt{(1-s_{ij})}$ we use $w_{ij}=1$ as edge weights, and we call this model Sparse Binary Graph (SBG). (II) We also consider a similarity threshold $th=0.4$ which yields a much denser graph, called Dense Weighted Graph (DWG). (iii) Finally, we compare these results with a “pose graph”, where instead of using DINOv2 features, we use distances between camera poses to build the affinity map. In this case we consider a binary graph where we cut edges for which the distance between the camera poses is above $1$m. We denote the Pose Graph with PG. From the results shown in Tab. 10 we can draw the following observations. First, using a pose aware graph (PG) does not necessarily improve the accuracy over using DINOv2-Graph, and when it does, the improvement is rather small. Note however that the pose of the observation and goal images were not used to select the start and end points of the shortest path, $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$, these were retrieved using DINOv2 feature similarity and we only replaced the DINOv2-Graph with the Pose Graph. The second observation we can make is that the graph structure does not have a big impact on the model with a Gumbel selector, but a bigger one on the model with attention-based context encoder. This is not surprising as the Gumbel selects an image from the context and use it independently from other context elements, while Attention combines information from all images within the context, hence it is more important how we fill up the context. Finally, we observe that the Dense Weighted Graph (DWG) performs better than the Sparse Weighted Graph (SWG). We conjecture that the reason is that the model was trained with $\mathcal{D}$ of size $10$k images and we rarely had no connection between two requested nodes ($r_{1}^{\mathbf{x}_{t}}$ , $r_{1}^{\mathbf{g}}$), which occurs more often at test when only $1$k images are available. Using a denser graph (DWG) probably helps overcoming these cases.
@@
Similarities and differences with topological maps – The proposed graph representation is related to topological map structures [63, 4, 14, 76, 40, 69]. However there are several critical differences between our model and traditional topological map based methods. First, we do not use the topological map to localize the agent, plan, and act to reach the “next waypoint”. Instead, the information gathered from planning is passed to the agent as a “recommendation” in the form of the retrieval context, to be integrated into decision making through learning: our agent is a full fledged policy which has access to memory, goal and a series of images along the path, which arguably enables more informed decision-making, and improved robustness and navigation performance. Additionally, our graph is connected based only on visual similarity rather than predictions from temporal distances. While it is debatable whether this approach is inherently superior, it offers a distinct advantage: we do not require trajectories to build the graph but can build it from a large number of isolated images. The main drawback is that we rely on the assumption that visual proximity is a proxy for spatial proximity, which is not always true (see Fig. 6). Future work could attempt to learn to take into account the uncertainty in this assumption through data-driven planning, as in [4].
Appendix B Additional ablation experiments
@@
In this section, we provide additional ablation studies that could not fit into the main document. As for the experiments in Sec. 5.4 of the main paper, all experiments are conducted on RANa-t models with Gumbel selector as context encoder $c$.
4 8 12 16 20 24
@@
Impact of context size – is studied in Tab. 11 for ImageNav ([Uncaptioned image]) and in Tab. 12 for ObjectNav ([Uncaptioned image]). One interesting advantage of the Gumbel selector is that it allows for context of different sizes during evaluation, independent of the context size used during training. With that said, this parameter has a relatively minor impact on performance, with larger training context sizes showing positive performance trends. The size $N=8$ is a choice that works well both for ImageNav and ObjectNav and balances navigation performance with execution time. For ObjectNav, we observe a more pronounced performance increase for larger context sizes during evaluation, so we use $N=8$ during training and $N=12$ during evaluation.
[…]
Model Context SR SPL
RANa-t top 8 83.0 52.3
RANa-t top 1 79.6 48.7
Figure 7: Refer to caption  Histogram of the rank of the context element selected by RANa-t with Gumbel context encoder on the ImageNav Gibson eval set.
Impact of retrieval database size – on zero-shot Instance-ImageNav ([Uncaptioned image]) is shown on Tab. 13. Here the goal image $\mathbf{g}^{\mathcal{R}}$ is retrieved from the database, and thus retrieval accuracy is crucial for success and avoid terminating an episode too far from the actual goal. Tab. 13 shows this phenomenon, with zero-shot Instance-ImageNav performance clearly increasing with retrieval dataset size.
1k 5k 10k 20k 50k 100k
47.6 (20.2) 56.2 (24.4) 56.4 (23.8) 57.4 (25.0) 58.7 (26.8) 59.1 (25.3)
@@
Analysis of “oracle” context – for ImageNav ([Uncaptioned image]) is carried out in Tab. 14. In this experiment we study which information is most useful and can be exploited by RANa, by building contexts using privileged information in simulation – therefore in these cases the retrieval dataset $\mathcal{D}$ is not used. Oracle 1 captures a panoramic view of the goal location and is created by collecting $N=8$ context images at the goal position while rotating the camera view by $45^{\circ}$ clock-wise for each element. This context represents a soft upper bound for the static context in Sec. 4.1 , where a fixed set of items is chosen per episode based on their similarity with the goal image. Oracle 2 context is richer and it is generated by computing the shortest path from the agent position to the goal at each step and collecting views along this path. This is a soft upper bound for the DINOv2-Graph context (Sec. 4.2 ), which is dynamically generated at each step on the context graph. Tab. 14 shows that, as expected, both types of oracle context can provide large gains, especially in SPL, i.e.navigation efficiency. The Table suggests that there is little room for improvement in success (SR) using static goal context, while more structured information, e.g.paths, derived from data in $\mathcal{D}$ can improve navigation considerably.
@@
RANa-t goal static 83.0 52.3
@@
RANa-t DINOv2-graph dynamic 84.7 63.7
@@
Table 14: [Uncaptioned image]  Navigation with “oracle” context - SR (SPL) for ImageNav using privileged information to build oracle context.
Impact of using pre-trained GRU weights – As mentioned earlier all RANa agents use a GRU finetuned from the baseline agent’s GRU. We evaluate the impact of this choice for the ImageNav ([Uncaptioned image]) model by training the RANa-t GRU from scratch while keeping all other settings fixed. Tab. 15 shows that finetuning the GRU form pretrained weights – those of DEBiT [6] in this case – improves navigation performance. We also observe faster convergence during training, which might in part explain the performance increase.
@@
Table 15: [Uncaptioned image]  Impact of GRU finetuning - SR (SPL) for ImageNav with GRU finetuned and trained from scratch.
Context selection criteria – While in RANa ObjectNav the context selection is always based on the observation, as the goal image is not available, in RANa ImageNav model we have both observation and goal images. All ImageNav experiments reported in the paper use a feature vector for each context element which is obtained by concatenating two features: (i) the output of the Geometric FM $d$ comparing context element and observation, and (ii) the output of $d$ comparing context element and goal. We evaluate the impact of this choice by training a RANa-t model which selects context elements based only on feature (i) which compares context elements to the observation. The performance of this variant, compared to the baseline which uses both observation and goal, is displayed in Tab. 16, which shows that information provided by comparing the context to the goal is useful and can be exploited by RANa to navigate more successfully and efficiently to the goal.
@@
Table 16: [Uncaptioned image]  Impact of context selection - SR (SPL) for ImageNav with context elements selected based on comparisons with only the observation, or both goal and observation.
Context building for ObjectNavTab. 17 shows the behavior of RANa-t ObjectNav ([Uncaptioned image]) when adopting different context building approaches, (i) rand: 8 random elements from the retrieval dataset, (ii) top 8 $cl_{S}$: most similar database images to the object goal encoder in CLIP feature space using CLIP similarity as metric, and (iii) top 8 $\omega_{\mathbf{o}}$: most similar database images using the score softmax normalization described in Sec. (4.1 ) of the main paper. The observed behavior is similar to ImageNav (Sec. 5.4 ). First, a RANa-t model trained with a context with randomly selected images from the scene gallery performs worse than a model trained with either top 8 images. Nevertheless, this model still outperforms our baseline trained without retrieval (SR=$42.9$ and SPL=$23.0$ for the non-retrieval baseline) and is more robust to random context images during evaluation. For RANa-t models trained with retrieval top 8 context, we observe the improved performance of using softmax normalized scores.
@@
Table 17: [Uncaptioned image]  Impact of retrieval data selection - SR (SPL) for ObjectNav with different context building strategies, combinations of train and test conditions.
Impact of retrieval dataset size for ObjectNav – is shown in Tab. 18. Results demonstrate that unlike ImageNav (Tab. 5 ), ObjectNav seems to need larger retrieval datasets both at training and at testing. Performance degrades significantly when the dataset $\mathcal{D}$ at training only has $1,000$ images. At test, using larger retrieval datasets leads to better performance, with metrics degrading significantly with $\mathcal{D}$ containing $100$ items.
100 1k 10k 50k
@@
Table 18: [Uncaptioned image]  Impact of retrieval database size - SR (SPL) for ObjectNav with retrieval using different database sizes, combinations of train and test conditions.
@@
Impact of CLIP variants for ObjectNav – is evaluated in Tab. 19. In this experiment, we build retrieval datasets $\mathcal{D}$ of size $10$k images for each of the 100 HM3DSem-v0.2 scenes used for ObjectNav (80 train, 20 test). For each of the 6 target object categories (chair, bed, plant, toilet, tv monitor, sofa) and for each scene, we do a search in the retrieval dataset $\mathcal{D}$ by querying the CLIP-encoded goal category (i.e.the CLIP text encoder for the sentence “a picture of a category inside a house”) and comparing it with CLIP encoded dataset images of the scene, providing top 1, top 8 and top 20 success percentages. A retrieval is considered successful if the retrieved image (whose ground-truth pose is known), is closer to one of the viewpoints of a object instance of the correct category than the success threshold of the ObjectNav task, i.e.1111m. The Table shows that OpenCLIP performs better than alternative CLIP models, namely OpenAI’s original CLIP weights [59] and SigLIP [85], all with the softmax normalization proposed in Sec. 4.1 being used. As observed earlier, ranking images by CLIP similarity to the category does not perform well [58]; since the number of target classes is known (6 in all our experiments), rescaling the score of each image in $\mathcal{D}$ via softmax significantly boosts performance.
Vision-language model softmax (Sec. 4.1 ) top 1 top 8 top 20
OpenCLIP [35] 40.6 76.0 88.5
OpenCLIP [35] 78.1 95.8 97.9
CLIP [59] 67.7 88.5 90.6
SigLIP [85] 70.8 91.7 94.8
@@
Table 19: Retrieval performance of CLIP variants in the context of ObjectNav ([Uncaptioned image]). Retrieval is considered success if the retrieved image is of the correct category and if it is closer than $1$m to the target object.
Appendix C Qualitative analysis
C.1 Examples of retrieved context
In this section we provide visual examples of the different context types used by our RANa model.
Fig. 8 shows three types of context used in this work for RANa ImageNav ([Uncaptioned image]), for $N=8$: top 8 most similar dataset images to the image goal in DINOv2 feature space (top) and top 8 MMR most similar images filtered by Maximum Margin Relevance (MMR) (middle), are static context types, i.e.they are fixed along one episode and depend only on the goal image, shown on the right. DINOv2-Graph (bottom) is a dynamic context that is recalculated at each step and depends on the observation (left) as well.
This example clearly shows that the most similar images to the goal are very similar to each other as well, and might not be very useful to the RANa agent. Filtering images using MMR increases the diversity of the context items, and this, as shown in Tab. 9 in the main text, increases RANa performance. For navigation purposes, the DINOv2-Graph context seems the most informative, as it provides waypoints from the current location to the goal – indeed this context provides the best navigation performance for RANa, see Tab. 2 .
Figure 8: Refer to caption  Context types for ImageNav - From top to bottom: top 8 most similar dataset images to the image goal (right), top 8 MMR most similar images filtered by MMR, and DINOv2-Graph, the only one depending on the observation (left) as well.

### v2

@@ Appendix C DINOv2-Graph
Selecting a path on this semantic graph (DINOv2-Graph) is not necessarily equivalent to a “physical path” (see examples in Fig.˜7). In order to enforce selection of edges between images that have high overlap probability (i.e. represent the same part of the scene) we made two further modifications to the above graph. First, as we observed that a DINOv2 similarity score above $th=0.75$ between L2 normalized features reflects, in most cases, a good overlap between the two images, we only kept edges for which the similarity was above this threshold, removing all the others. Second, in order to enforce Dijkstra’s algorithm to select paths with high similarity (low cost), we used the weights $w_{ij}=\sqrt{(1-s_{ij})}$, where $s_{ij}$ is the DINOv2 similarity between the L2 normalized features.
During both training and inference, at each step, the agent finds in $\mathcal{D}$ the most similar images $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ to the observation $\mathbf{x}_{t}$ and goal $\mathbf{g}$, respectively, and computes the shortest path between $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$ on the DINOv2-Graph. Let $\mathcal{P}=\{r_{1}^{\mathbf{x}_{t}},r_{1},r_{2},...,r_{n},r_{1}^{\mathbf{g}}\}$ be the set of nodes on the shortest path and $C$ the context size. If $n+2>C$, we sample images randomly from $\mathcal{P}$ and update all images in the context. If $n+2\leq C$, we use all images in $\mathcal{P}$ to update $n+2$ images in the context. If no path was found, $n=2$ hence we only update two elements of the context with $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$.
Analyzing the choice of the graph structure – We trained our models on the Gibson dataset where we had around 101010k images per scene in the retrieval dataset $\mathcal{D}$. We selected the checkpoint obtained at 130130130M steps for all models. During testing we used a set of $|\mathcal{D}|=1$k images per scene. In this section we study how the graph structure affects performance. The default model, used for the results reported in Table˜2 of the main paper, builds a weighted graph by removing edges with similarity below $th=0.75$ and uses $w_{ij}=\sqrt{(1-s_{ij})}$ as weights. We call this Sparse Weighted Graph (SWG). We evaluate three additional strategies to construct the DINOv2-Graph. (i) We consider the same graph, but instead of using $w_{ij}=\sqrt{(1-s_{ij})}$ we use $w_{ij}=1$ as edge weights, and we call this model Sparse Binary Graph (SBG). (ii) We also consider a similarity threshold $th=0.4$ which yields a much denser graph, called Dense Weighted Graph (DWG). (iii) Finally, we compare these results with a “pose graph”, where instead of using DINOv2 features, we use distances between camera poses to build the affinity map. In this case we consider a binary graph where we cut edges for which the distance between the camera poses is above 111m. We denote the Pose Graph with PG.
@@
Table 10: [Uncaptioned image]  Impact of the graph structure on RANa with DINOv2-Graph - SR (SPL) for ImageNav with retrieval using different graph structures at test time: Sparse Weighted Graph (SWG), Sparse Binary Graph (SBG), Dense Weighted Graph (DWG) and Pose Graph (PG). All RANa models were trained with the same SWG graphs built on the training scenes. In bolded best results. Note that the PG variant is not considered as it uses the camera poses of images in $\mathcal{D}$; we only add it for comparison.
@@
From the results shown in Table˜10 we can draw the following observations. First, using a pose aware graph (PG) does not necessarily improve the accuracy over using DINOv2-Graph, and when it does, the improvement is rather small. Note however that the pose of the observation and goal images were not used to select the start and end points of the shortest path, $r_{1}^{\mathbf{x}_{t}}$ and $r_{1}^{\mathbf{g}}$, these were retrieved using DINOv2 feature similarity and we only replaced the DINOv2-Graph with the Pose Graph. The second observation we can make is that the graph structure does not have a big impact on the model with a Gumbel selector, but a bigger one on the model with attention-based context encoder. This is not surprising as the Gumbel selects an image from the context (path) and use it independently from other context elements, while Attention combines information from all images within the context (path), hence it is more important how we fill up the context. Finally, we observe that the Dense Weighted Graph (DWG) performs better than the Sparse Weighted Graph (SWG). We conjecture that the reason is that the model was trained with $\mathcal{D}$ of size 101010k images and we rarely had no connection between two requested nodes ($r_{1}^{\mathbf{x}_{t}}$ , $r_{1}^{\mathbf{g}}$), which occurs more often at test when only 111k images are available. Using a denser graph (DWG) appears to help overcoming these cases.
@@
Similarities and differences with topological maps – The proposed graph representation is related to topological map structures (Savinov et al., 2018; Beeching et al., 2020b; Chaplot et al., 2020c; Wiyatno et al., 2022; Kim et al., 2023; Sridhar et al., 2024). However, there are several critical differences between our model and traditional topological map based methods. First, we do not use the topological map to localize the agent, plan, and act to reach the “next waypoint”. Instead, the information gathered from planning is passed to the agent as a “recommendation” in the form of the retrieval context, to be integrated into decision making through learning: our agent is a full fledged policy which has access to memory, goal and a series of images along the path, which arguably enables more informed decision-making, and improved robustness and navigation performance. Additionally, our graph is connected based only on visual similarity rather than predictions from temporal distances. While it is debatable whether this approach is inherently superior, it offers a distinct advantage: we do not require trajectories to build the graph but we can also build it from a large number of isolated images. The main drawback of our method is that we rely on the assumption that visual proximity is a proxy for spatial proximity, which is not always true (see Fig.˜7). Future work could attempt to learn to take into account the uncertainty in this assumption through data-driven planning, as in Beeching et al. (2020b).
Appendix D Ablation experiments
@@
In this section, we provide additional ablation studies that could not fit into the main document. As for the ablation experiments in the main paper, all experiments are conducted on RANa-t models.
Impact of context size – is studied in Table˜11 for ImageNav ([Uncaptioned image]) and in Table˜12 for ObjectNav ([Uncaptioned image]) with static context. One interesting advantage of the Gumbel selector is that it allows for context of different sizes during evaluation, independent of the context size used during training. With that said, this parameter has a relatively minor impact on performance, with larger training context sizes showing positive performance trends. The size $N=8$ is a choice that works well both for ImageNav and ObjectNav and balances navigation performance with execution time. For ObjectNav, we observe a more pronounced performance increase for larger context sizes during evaluation, so we use $N=8$ during training and $N=12$ during evaluation.
context size 4 8 12 16 20 24
@@
Impact of context diversity for Gumbel context encoderTable˜13 evaluates the behavior of RANa-t ImageNav with static context, when using different strategies to create the context, (i) top 8 MMR: most similar database images filtered by MMR as in Section˜4.1, (ii) top 8: most similar database images to the goal in DINOv2 feature space, and (iii) top 1: the closest element from the retrieval dataset.
Filtering the context with MMR is important, as the agent appears to rely on diverse context items; removing MMR degrades performance. top 1 always uses the best element from the context, which is very similar to the goal image. In this case performance go back to baseline levels, confirming the importance of diversity in the retrieved elements.
top 8 MMR top 8 top 1
top 8 MMR 83.0 (52.3) 81.5 (51.0) 79.6 (48.7)
top 8 80.7 (51.1) 81.4 (48.1) 80.2 (51.2)
Table 13: Impact of MMR filtering on retrieved context - SR (SPL) for ImageNav with and without MMR, combinations of train and test conditions ( in-domain performance shaded in pink). Context diversity from MMR is important and delivers the best performance (SR=83.0). When the model trained with MMR is tested without it, performance drops.
Impact of retrieval database size – on zero-shot Instance-ImageNav ([Uncaptioned image]) is shown on Table˜14. Here the goal image $\mathbf{g}^{\mathcal{R}}$ is retrieved from the database, and thus retrieval accuracy is crucial for success and avoid terminating an episode too far from the actual goal. Table˜14 shows this phenomenon, with zero-shot Instance-ImageNav performance increasing with retrieval dataset size. However, please note that RANa achieves navigation performance in line with SOTA using a retrieval dataset containing only 5,0005,0005 , 000 images.
DB size 1k 5k 10k 20k 50k 100k
SR (SPL) 47.6 (20.2) 56.2 (24.4) 56.4 (23.8) 57.4 (25.0) 58.7 (26.8) 59.1 (25.3)
@@
Analysis of “oracle” context – for ImageNav ([Uncaptioned image]) is carried out in Table˜15. In this experiment we study which information is most useful and can be exploited by RANa, by building contexts using privileged information in simulation – therefore in these cases the retrieval dataset $\mathcal{D}$ is not used. Oracle 1 captures a panoramic view of the goal location and is created by collecting $N=8$ context images at the goal position while rotating the camera view by $45^{\circ}$ clock-wise for each element. This context represents a soft upper bound for the static context in Section˜4.1, where a fixed set of items is chosen per episode based on their similarity with the goal image. Oracle 2 context is richer and it is generated by computing the shortest path from the agent position to the goal at each step and collecting views along this path. This is a soft upper bound for the DINOv2-Graph dynamic context (Section˜4.2), which is dynamically generated at each step on the context graph. Table˜15 shows that, as expected, both types of oracle context can provide large gains, especially in SPL, i.e. navigation efficiency. It suggests that there is little room for improvement in success (SR) using static goal context, while more structured information, e.g. paths, derived from data in $\mathcal{D}$ can improve navigation considerably.
@@
DINOv2 goal static 83.0 52.3
@@
DINOv2-graph dynamic 84.7 63.7
@@
Table 15: [Uncaptioned image]  Navigation with “oracle” context - SR (SPL) for RANa-t using privileged information to build oracle context.
Impact of using pre-trained GRU weights – As mentioned earlier all RANa agents use a GRU finetuned from the baseline agent’s GRU. We evaluate the impact of this choice by training the RANa-t static GRU from scratch while keeping all other settings fixed. Table˜16 shows that finetuning the GRU form pretrained weights – those of DEBiT (Bono et al., 2024a) in this case – improves navigation performance. We also observe faster convergence during training, which might in part explain the performance increase.
@@
Table 16: [Uncaptioned image]  Impact of GRU finetuning - SR (SPL) for RANa-t with static context and GRU finetuned or trained from scratch.
Context selection criteria – All ImageNav experiments reported in the paper use a feature vector for each context element which is obtained by concatenating two features: (i) the output of the Geometric FM $g$ comparing context element and observation, and (ii) the output of $g$ comparing context element and goal. We evaluate the impact of this choice by training a RANa-t static model which selects context elements based only on feature (i), which compares context elements to the observation. The performance of this variant and that of the baseline which uses both observation and goal, is displayed in Table˜17. It shows that information provided by comparing the context to the goal is useful and can be exploited by RANa to navigate more successfully and efficiently to the goal.
@@
Table 17: [Uncaptioned image]  Impact of context selection - SR (SPL) for RANa-t with static context and elements selected based on comparisons with only the observation, or both goal and observation.
Context building for ObjectNavTable˜18 shows the behavior of RANa-t ObjectNav ([Uncaptioned image]) when adopting different context building approaches, (i) rand: 8 random elements from the retrieval dataset, (ii) top 8 $cl_{S}$: most similar database images to the object goal encoder in CLIP feature space using CLIP similarity as metric, and (iii) top 8 $\omega_{\mathbf{o}}$: most similar database images using CLIP score softmax normalization. While CLIP allows to easily compare any open-vocabulary object category with retrieval gallery images, it has been observed in the context of zero-shot image classification that directly ranking images by similarity to the category is suboptimal (Qian et al., 2023). To address this issue for ObjectNav, we leverage the fact that we often target a pre-defined set of classes $\mathcal{O}$ (e.g. 6 in HM3D), and rescale the score of each image in $\mathcal{D}$ via softmax:
The observed behavior is similar to ImageNav. First, a RANa-t model trained with a context with randomly selected images from the scene gallery performs worse than a model trained with either top 8 images. Nevertheless, this model still outperforms our baseline trained without retrieval (SR=42.942.942.9 and SPL=23.023.023.0 for the non-retrieval baseline) and is more robust to random context images during evaluation. For RANa-t models trained with retrieval top 8 context, we observe improved performance when using softmax normalized scores.
@@
Table 18: [Uncaptioned image]  Impact of retrieval data selection - SR (SPL) for RANa-t ObjectNav with different context building strategies, combinations of train and test conditions.
Impact of retrieval dataset size for ObjectNav – is shown in Table˜19. Results demonstrate that unlike ImageNav (Table˜5), ObjectNav seems to need larger retrieval datasets both at training and at testing. Performance degrades significantly when the dataset $\mathcal{D}$ at training only has 1,0001,0001 , 000 images. At test, using larger retrieval datasets leads to better performance, with metrics degrading significantly with $\mathcal{D}$ containing 100100100 items.
DB size 100 1k 10k 50k
@@
Table 19: [Uncaptioned image]  Impact of retrieval database size - SR (SPL) for RANa-t ObjectNav with retrieval using different database sizes, combinations of train and test conditions.
@@
Impact of CLIP variants for ObjectNav – is evaluated in Table˜20. In this experiment, we build retrieval datasets $\mathcal{D}$ of size 101010k images for each of the 100 HM3DSem-v0.2 scenes used for ObjectNav (80 train, 20 test). For each of the 6 target object categories (chair, bed, plant, toilet, tv monitor, sofa) and for each scene, we do a search in the retrieval dataset $\mathcal{D}$ by querying the CLIP-encoded goal category (i.e. the CLIP text encoder for the sentence “a picture of a category inside a house”) and comparing it with CLIP encoded dataset images of the scene, providing top 1, top 8 and top 20 success percentages. A retrieval is considered successful if the retrieved image (whose ground-truth pose is known), is closer to one of the viewpoints of a object instance of the correct category than the success threshold of the ObjectNav task, i.e. 111m. The Table shows that OpenCLIP performs better than alternative CLIP models, namely OpenAI’s original CLIP weights (Radford et al., 2021) and SigLIP (Zhai et al., 2023), all with the softmax normalization being used. As observed earlier, ranking images by CLIP similarity to the category does not perform well (Qian et al., 2023); since the number of target classes is known (6 in all our experiments), rescaling the score of each image in $\mathcal{D}$ via softmax significantly boosts performance.
Vision-language model softmax top 1 top 8 top 20
OpenCLIP (Ilharco et al., 2021) 40.6 76.0 88.5
OpenCLIP (Ilharco et al., 2021) 78.1 95.8 97.9
CLIP (Radford et al., 2021) 67.7 88.5 90.6
SigLIP (Zhai et al., 2023) 70.8 91.7 94.8
@@
Table 20: Retrieval performance of CLIP variants. Retrieval is considered success if the retrieved image is of the correct category and if it is closer than 111m to the target object.
Appendix E Qualitative analysis
E.1 Examples of retrieved context for ObjectNav